Calculate the standard deviation of the following data;

45, 50, 67, 68, 75, 80, 70, 65, 70,90
1
Expert's answer
2021-06-28T03:53:25-0400

"n=10"


"\mu=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{10}(45+50+67+68+75+80"

"+70+65+70+90)=68"

"\sigma^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\mu)^2}{n}=\dfrac{1}{10}((45-68)^2+(50-68)^2"

"+(67-68)^2+(68-68)^2+(75-68)^2"

"+(80-68)^2+(70-68)^2+(65-68)^2"


"+(70-68)^2+(90-68)^2)=154.8"

"\sigma=\sqrt{\sigma^2}=\sqrt{154.8}\approx12.4419"